OpenStudy (anonymous):

*** I will give you a medal if you can help/answer *** Simplify the complex fraction. (I'm going to draw it) @jim_thompson5910

OpenStudy (tkhunny):

Have you considered the Complex Conjugate?

OpenStudy (anonymous):



OpenStudy (tkhunny):

Ph, you mean the realatively complicated fraction. i DOwish textbooks would quit caling these "Complex" fractions. VERY confusing! Is it this? \(\dfrac{x+ \dfrac{4x}{y}}{\dfrac{7}{3y}}\)

OpenStudy (unklerhaukus):

\[\frac{a}{b/c}=\frac{ac}b\]

OpenStudy (anonymous):

the 3y is 3x .. but yes that is what it looks like.

OpenStudy (tkhunny):

Justaket a piece at a time. \(x + \dfrac{4x}{y} = \dfrac{xy}{y} + \dfrac{4x}{y} =\dfrac{xy + 4x}{y}\)

OpenStudy (anonymous):

what about the 3 ?

OpenStudy (tkhunny):

You tell me. That's the next part. We have \(\dfrac{\dfrac{xy+4x}{y}}{\dfrac{7}{3x}}\)

OpenStudy (anonymous):

am i going to multiply the top part (what you just did) but the recip of the bottom part which would be (3x)/(7)

OpenStudy (tkhunny):

Perfect. That is exactly correct.

OpenStudy (anonymous):

okayy.. so the denominator would be 7y. (Still working on the problem)

OpenStudy (tkhunny):

Yes.

OpenStudy (anonymous):

and the numerator would be ... 3x^2(y+4) ?

OpenStudy (tkhunny):

That's it!

OpenStudy (anonymous):

that's the answer ? I don't distribute or anything? that's my final answer?

OpenStudy (tkhunny):

An important of mathmatical confidence is knowing whe you are done. Move on to the next one!

OpenStudy (anonymous):

Thank you (:

OpenStudy (tkhunny):

* important part